OpenStudy (anonymous):

Simplify. 9 – (3)(–5) A. –30 B. –6 C. 30 D. 24

OpenStudy (alexandervonhumboldt2):

3*(-5)=-(3*5)=????

OpenStudy (alexandervonhumboldt2):

so we have 9-(-(3*5))=9-(-15)=9+15=???

OpenStudy (jadedry):

In an equation you have to multiply/divide before you subtract/add. The exception to this rule is when subtraction/addition expressions are bracketed.